24. Amazing Gizmos, Inc. originally forecasted the following financial data for next year: Sales = $874,000, Cost of goods sold = $589,950, operating expense and depreciation = $125,675 and interest expense = $39,875. The firm believes that COGS will always be 67.5% of sales. Suppose the firm wants to achieve a net income of $112,500. Assuming the operating costs, depreciation, and interest expense will remain the same, how large must sales be to achieve this goal? Assume a 28.0% tax rate.


25. The current balance sheet for Tampa Tanks, Inc. lists net fixed assets as $736,000. The fixed assets could currently be sold for $912,000. The balance sheet shows current liabilities of $328,575 and net working capital of $143,860. If all the current accounts were liquidated today, the company would receive $216,400 cash after paying all of its current liabilities at book value. How much larger is market value than book value of the firm’s assets today?
Answer #1

(24) Target Net Income = $ 112500, Tax Rate = 28 %

Profit Before Tax = 112500 / (1-Tax Rate) = 112500 / (1-0.28) = $ 156250

Add: Interest Expense = $ 39875

EBIT = $ 196125

Add: Operating Costs and Depreciation = $ 125675

Gross Profit = $ 321800

As Cost of Goods Sold is always 67.5 % of Sales, then assuming sales is $ Y, the Gross Profit = Y - 0.675Y

Therefore, 0.325Y = 321800

Y = 321800 / 0.325 = $ 990153.85

NOTE: Please raise a separate query for the solution to the second unrelated question, as one query is restricted to the solution of only one complete question with up to four sub-parts.